OpenStudy (anonymous):

If h(2)=9 and h'(2)=-2, find the following

OpenStudy (anonymous):

\[\frac{ d }{ dx }(\frac{ h(x) }{ x })\]

OpenStudy (anonymous):

X=2

OpenStudy (anonymous):

so what i thought is that since h(2)=9, they want the derivative of 9/2 which is zero

OpenStudy (anonymous):

apparently thats not the case

OpenStudy (anonymous):

use the quotient rule: \[\frac{ d }{ dx }\left( \frac{ f(x) }{ g(x) } \right)= \frac{ f'(x)g(x) - f(x)g'(x) }{ (g'(x))^2 }\]

OpenStudy (anonymous):

I also did that and didnt check my work lol i thought i divided by zero

OpenStudy (anonymous):

hehe lol